# ReTool: Reinforcement Learning for Strategic Tool Use in LLMs

### v1

@@ 1 Introduction
We evaluate ReTool on the challenging MATH Olympiad benchmarks AIME2024 and AIME2025. Building on Qwen2.5-32B-Instruct [29], our model achieves 67.0% accuracy on AIME2024 with only 400 training steps, significantly outperforming the text-based RL baseline, which achieves 40.0% accuracy with 1080 training steps. These substantial gains highlight that explicitly modeling tool-use as part of the decision process not only pushes the limits of model reasoning but also enhances training efficiency. Furthermore, when trained on DeepSeek-R1-Distill-Qwen-32B [4], our model demonstrates further improvements, surpassing competitive baselines such as QwQ-32B-Preview [23], s1-32B [10], and OpenAI o1-preview [11]. This suggests that the RL training process inspires more efficient problem-solving strategies. Additionally, our cold-start model based on Qwen2.5-32B-Instruct achieves 40.9% accuracy on AIME2024, demonstrating performance on par with the text-based RL baseline (40.0%). This indicates that the cold-start dataset constructed via our automated pipeline effectively captures tool usage patterns within executable reasoning traces, providing an initialization for downstream RL training. We further conduct a comprehensive analysis of CI cognitive behavior through RL training and identify several key findings. Our model demonstrates enhanced code utilization capabilities, enabling it to employ more accurate and complex code snippets; It also learns to invoke tools appropriately, select tool adaptively, structure tool calls effectively, and iteratively refine reasoning through emergent code self-correction capabilities.
@@ 2 Methodology
In this section, we introduce ReTool, a CI-powered RL framework designed to address math problem-solving tasks. We begin with an overview of ReTool. Next, we describe our cold start training, including the data construction pipeline and supervised fine-tuning (section 2.2). We then outline our reinforcement learning pipeline, enhanced by a code interpreter sandbox, to further enhance strategic tool usage development (section 2.3).
@@ 2.1 Overview
Our methodology consists of two primary stages: cold start supervised fine-tuning followed by reinforcement learning with interleaved code execution rollout. Firstly, we collect data through our designed pipeline for cold start supervised fine-tuning (SFT), which provides a robust initialization for the reinforcement learning phase. To enhance our model’s tool utilization capabilities, we introduce a specialized tool-using reinforcement learning pipeline that enhances the model’s ability to appropriately select and apply tools during the reasoning process.
2.2 Cold Start for Tool-Integrated Reasoning Foundation
@@
ReTool employs SFT to learn when and how to invoke the code interpreter from the aforementioned dataset $\mathcal{D}_{\text{CI}}$, thereby enhancing the model’s capability to appropriately utilize computational tools.
@@ 2.3.2 Training Details
KV-Cache Reuse. In order to reduce the memory cost during rollout, when each time the is code termination trigger (</code>) is detected, we will cache all the KV-cache before code execution and only calculate and append the KV-cache from the interpreter feedback (<interpreter></interpreter>). This will largely reduce the KV-cache for each rollout.
@@ 3 Experiment
3.1 Experimental Setup
For training, we employ the VeRL framework [^1] https://github.com/volcengine/verl. We adopt PPO as our RL method. Regarding hyperparameters, we utilize the AdamW optimizer with an initial learning rate of 1e-6. We define the expected maximum sequence length as 16384 tokens. For training, the mini-batch size is set to 512, and the KL coefficient is set to 0.0. We use Qwen2.5-32B-Instruct [15] as the main backbone.
@@ 3.1 Experimental Setup
Ablations
Qwen2.5-32B-Instruct 26.7 -
Qwen2.5-32B-Instruct-ColdStart 40.9 34.5
Text-based RL (Qwen2.5-32B-Instruct) 40.0 36.7
Table 1: Main results.
@@ 3.2 Main Results
As shown in Table 1, ReTool enables the LLM to flexibly utilize the code interpreter during the RL stage, resulting in notable improvements in performance. Specifically, ReTool (Qwen2.5-32B-Instruct) achieves scores of 67.0% and 49.3% on AIME2024 and AIME2025, respectively, with only 400 training steps. This performance significantly surpasses the text-based RL baseline (Qwen2.5-32B-Instruct), which attains 40.0% and 36.7% with over 1000 training steps. These results suggest that the tool-integrated learning approach of ReTool not only extends the reasoning capabilities of the model but also enhances training efficiency. On AIME2024, ReTool (Qwen2.5-32B-Instruct) outperforms the competitive baseline s1-32B by 10.3%. Similarly, on AIME2025, ReTool (Qwen2.5-32B-Instruct) achieves a 11.4% improvement over OpenAI o1-preview. This demonstrates that more efficient problem-solving strategies have been discovered during RL training process. Furthermore, when paired with a more advanced backbone model, ReTool (DeepSeek-R1-Distill-Qwen-32B) further enhances performance, achieving scores of 72.5% and 54.3% on AIME2024 and AIME2025, respectively. These results collectively demonstrate the effectiveness of ReTool.
@@ 3.3 Cognitive Analysis
“Aha Moment” of Code Self-correction. We surprisingly find that our model demonstrates the ability to self-correct non-executable code, despite the absence of explicit training data for code self-correction. As illustrated in Figure 4, the model initially generated code that failed to execute due to a missing ‘NumPy’ import. Based on interpreter feedback, the model identified the flaw by generating the reflection, “Oops, forgot to import numpy. Let’s correct that”, followed by an executable version of the code that included the necessary import statement. This emergent behavior suggests that reinforcement learning can foster metacognitive capabilities, enabling the model to iteratively refine its generated code to address more complex problems.
@@ 4.2 Tool Integrated Reasoning
Tool-integrated reasoning was first introduced to help LLMs solve computationally intensive mathematical problems with the integration of programming strategies [1, 32, 6, 18]. Building on this foundation, Wang et al. [24] proposed an iterative approach that combines textual reasoning with code execution to mutually verify and enhance reasoning accuracy. More recently, Chen et al. [2] integrated code execution into the reasoning process by performing supervised fine-tuning on self-curated code-integrated CoT data. However, this approach is inherently limited by its reliance on the specific data distribution, and cannot learn adaptive strategies for tool use—such as determining when and how to invoke tools—through reinforcement learning. A concurrent work [8] applied reinforcement learning to learn tool usage strategies on Qwen2.5-Math models [30] at 1.5B and 7B scales, but the performance remained suboptimal. We further scale up this line of research and propose ReTool, a framework that leverages reinforcement learning to strategically determine when and how to invoke the code interpreter. Our method outperforms Qwen-Math-72B-TIR [30] and o1-preview [11] significantly on AIME2024 and AIME2025. We also present a comprehensive analysis of the learned tool-use behaviors and highlight several key findings regarding the model’s cognitive patterns in code invocation after ReTool training.
@@ Acknowledgments
We thank other colleagues at ByteDance for their support for the ReTool project. The corresponding author is Wanjun Zhong (wanjun@bytedance.com).

### v2

@@ 1 Introduction
We evaluate ReTool on the challenging MATH Olympiad benchmarks AIME2024 and AIME2025. Building on Qwen2.5-32B-Instruct [30], our model achieves 67.0% accuracy on AIME2024 with only 400 training steps, significantly outperforming the text-based RL baseline, which achieves 40.0% accuracy with 1080 training steps. These substantial gains highlight that explicitly modeling tool-use as part of the decision process not only pushes the limits of model reasoning but also enhances training efficiency. Furthermore, when trained on DeepSeek-R1-Distill-Qwen-32B [4], our model demonstrates further improvements, surpassing competitive baselines such as QwQ-32B-Preview [23], s1-32B [10], and OpenAI o1-preview [11]. This suggests that the RL training process inspires more efficient problem-solving strategies. Additionally, our cold-start model based on Qwen2.5-32B-Instruct achieves an accuracy of 40.9% on AIME2024, comparable to the text-based RL baseline based on same backbone (40.0%), and significantly surpasses the non-trained Qwen2.5-32B-Instruct (26.7%). These results demonstrate that our curated dataset effectively captures tool usage patterns within executable reasoning traces, and that CI-integrated training positively contributes to reasoning performance.
We further conduct a comprehensive analysis of CI cognitive behavior through RL training and identify several key findings. Our model demonstrates enhanced code utilization capabilities, enabling it to employ more accurate and complex code snippets; It also learns to invoke tools appropriately, select tool adaptively, structure tool calls effectively, and iteratively refine reasoning through emergent code self-correction capabilities.
@@ 2 Methodology
In this section, we introduce ReTool, a CI-powered RL framework designed to address math problem-solving tasks. We begin with an overview of ReTool. Next, we describe our cold-start training, including the data construction pipeline and supervised fine-tuning (section 2.2). We then outline our reinforcement learning pipeline, enhanced by a code interpreter sandbox, to further enhance strategic tool usage development (section 2.3).
@@ 2.1 Overview
Our methodology consists of two primary stages: cold-start supervised fine-tuning followed by reinforcement learning with interleaved code execution rollout. Firstly, we collect data through our designed pipeline for cold-start supervised fine-tuning (SFT), which provides a robust initialization for the reinforcement learning phase. To enhance our model’s tool utilization capabilities, we introduce a specialized tool-using reinforcement learning pipeline that enhances the model’s ability to appropriately select and apply tools during the reasoning process.
2.2 Cold-start for Tool-Integrated Reasoning Foundation
@@
ReTool employs supervised fine-tuning to learn when and how to invoke the code interpreter from the aforementioned dataset $\mathcal{D}_{\text{CI}}$, thereby enhancing the model’s capability to appropriately utilize computational tools.
@@ 2.3.2 Training Details
Cold-start & RL For training, we employ the VeRL framework [^1] https://github.com/volcengine/verl. We adopt PPO as our RL method. We train our model on curated cold-start data for two epochs. Regarding hyperparameters, we utilize the AdamW optimizer with an initial learning rate of 1e-6. We define the expected maximum sequence length as 16384 tokens. For training, the mini-batch size is set to 512, and the KL coefficient is set to 0.0. We use Qwen2.5-32B-Instruct [15] as the main backbone.
@@
KV-Cache Reuse. In order to reduce the memory cost during rollout, when each time the code termination trigger (</code>) is detected, we will cache all the KV-cache before code execution and only calculate and append the KV-cache from the interpreter feedback (<interpreter></interpreter>). This will largely reduce the KV-cache for each rollout.
@@ 3 Experiment
3.1 Evaluation Setup
@@
Ablations on Qwen2.5-32B-Instruct
w/o Training (Base Model) 26.7 -
w/o CI (Text-based RL) 40.0 36.7
w/o RL (only Cold-start) 40.9 34.5
Table 1: Main results. The Text-based RL method includes a text-based cold-start SFT initialization to ensure a fair comparison. The inference process of the Cold-start model also incorporates code interpreter.
@@ 3.2 Main Results
As shown in Table 1, ReTool enables the LLM to flexibly leverage the code interpreter during the RL stage, leading to substantial performance improvements. Specifically, ReTool (Qwen2.5-32B-Instruct) achieves accuracies of 67.0% on AIME2024 and 49.3% on AIME2025 with only 400 training steps. This markedly outperforms the text-based RL baseline (Qwen2.5-32B-Instruct), which attains 40.0% and 36.7% on the respective benchmarks despite using over 1000 training steps. These findings indicate that the tool-integrated learning paradigm employed by ReTool not only enhances the model’s reasoning capabilities but also improves training efficiency. Furthermore, on AIME2024, ReTool (Qwen2.5-32B-Instruct) surpasses the competitive baseline s1-32B by 10.3%. Similarly, on AIME2025, it achieves an 11.4% gain over OpenAI’s o1-preview. When combined with a more advanced backbone, ReTool (DeepSeek-R1-Distill-Qwen-32B) further improves performance, achieving scores of 72.5% on AIME2024 and 54.3% on AIME2025. These results suggest that more effective problem-solving strategies are discovered during the RL training process.
Moreover, our cold-start model based on Qwen2.5-32B-Instruct achieves an accuracy of 40.9% on AIME2024, closely aligning with the performance of the text-based RL baseline (40.0%), and substantially surpassing the base model (26.7%). These results demonstrate that our curated dataset effectively captures tool usage patterns within executable reasoning traces, and that CI-integrated training contributes positively to reasoning performance.
@@ 3.3 Cognitive Analysis
“Aha Moment” of Code Self-correction. Interestingly, our model exhibits an emergent ability to self-correct non-executable code, despite the absence of explicit training data for code self-correction. As shown in Figure 4, the model initially produced code that failed to execute due to the undefined function “greedy()”. Upon receiving feedback from the interpreter, the model recognized the error and responded with the reflection: “Oops, the functions need to be defined in the same scope. Let’s correct that.” It then proceeded to generate a revised, executable version of the code that included all necessary function definitions. This emergent behavior suggests that reinforcement learning can foster metacognitive capabilities, enabling the model to iteratively refine its generated code to address more complex problems.
@@ 4.2 Tool Integrated Reasoning
Tool-integrated reasoning was first introduced to help LLMs solve computationally intensive mathematical problems with the integration of programming strategies [1, 33, 6, 18, 24]. Building on this foundation, Wang et al. [25] proposed an iterative approach that combines textual reasoning with code execution to mutually verify and enhance reasoning accuracy. More recently, Chen et al. [2] integrated code execution into the reasoning process by performing supervised fine-tuning on self-curated code-integrated CoT data. However, this approach is inherently limited by its reliance on the specific data distribution, and cannot learn adaptive strategies for tool use—such as determining when and how to invoke tools—through reinforcement learning. A concurrent work [8] applied reinforcement learning to learn tool usage strategies on Qwen2.5-Math models [31] at 1.5B and 7B scales, but the performance remained suboptimal. We further scale up this line of research and propose ReTool, a framework that leverages reinforcement learning to strategically determine when and how to invoke the code interpreter. Our method outperforms Qwen-Math-72B-TIR [31] and o1-preview [11] significantly on AIME2024 and AIME2025. We also present a comprehensive analysis of the learned tool-use behaviors and highlight several key findings regarding the model’s cognitive patterns in code invocation after ReTool training.
@@ Acknowledgments
We would like to thank Guang Shi, Mingxuan Wang, Renjie Zheng, Chen Dun, and Yun Jiang for their support on this work.